OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3. (4 points) @Directrix Could you help ill give metal out

OpenStudy (anonymous):

and become fan @Directrix

Directrix (directrix):

y = 2-x and y = 4x + 3 First, find the x-coordinate of the point of intersection of the lines with equations y = 2-x and y = 4x + 3. Then, show that the values of x makes 2-x = 4x + 3 true.

OpenStudy (anonymous):

hey is it possibel if you could help me with part B and C?

Directrix (directrix):

Where do these lines intersect: y = 2-x and y = 4x + 3 y = 2-x y = 4x + 3 ======== Substitute 2 - x for y in the second equation. 2-x = 4x + 3 ==> @rosiee123 --> See if you can solve this equation for x, Okay? Post what you get.

Directrix (directrix):

You are solving this for x: 2-x = 4x + 3

OpenStudy (anonymous):

i really dont know how to do it:P

Directrix (directrix):

2 -x = 4x + 3 -2 -2 ================ 0 - x = 4x + 1 ---> Do you agree with subtracting 2 from both sides?

Directrix (directrix):

@rosiee123

Directrix (directrix):

0 - x = 4x + 1 -4x -4x ---------------- -5x = 1 x = ? @rosiee123

Directrix (directrix):

@rosiee123 You must post in the thread and respond to questions. It is okay if you post wrong answers but you must participate or I cannot help you.

Directrix (directrix):

Let me complete this and maybe you can use it as a pattern for other problems like it.

Directrix (directrix):

-5x = 1 so x = -1/5

Directrix (directrix):

So , now show that x = -1/5 is a solution of the equation: 2-x = 4x + 3. ------------------------------------------------------------------- Strategy: In 2-x = 4x + 3, replace x by - 1/5 and show that both sides of the equation are equal. 2-x 4x + 3 2 - (-1/5) 4(-1/5) + 3 2 + 1/5 -4/5 + 3 2 1/5 = 2 1/5 The two sides of the equation are both equal to the mixed number 2 1/5 which can be written as 11/5. So, the x-coordinate of the solution of the system of equations is the value of x that makes the equation true. @rosiee123